The perimeter of a rectangle is 124cm.If the length is 12 more than the width, find the length and the width.

Answer by TimothyLamb(4379)

---
x = length
y = width
---
2x + 2y = 124
x = y + 12
---
put the system of linear equations into standard form:
---
2x + 2y = 124
x - y = 12
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length = 37 cm
y = width = 25 cm
---
system matrix
2 2
1 -1
---
inverse of system matrix
0.25 0.5
0.25 -0.5
---
determinant of system matrix = -4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php